OpenStudy (anonymous):

Given: h(x) = (1)/(2x + 3) ; g(x) = (x +1)/(x - 3) 1. (g of h)(x) 2. (h of g)((squareroot of 3) / 2) 3. 2/3 [(g of h)((squareroot of 2)/ 2)

OpenStudy (jhannybean):

Wait a minute.... is it (1)/(2x+3)???

OpenStudy (zzr0ck3r):

g(x) says x-3

OpenStudy (jhannybean):

\[\large h(x) =\frac{1}{2x+3}\]\[\large g(x)= \frac{x+1}{x-3}\]1. \[\large g(h(x)) = \frac{\frac{1}{(2x+3)}+1}{\frac{1}{2x+3}-3}\]

OpenStudy (jhannybean):

Omg person's offline. It's too much to type up :|

OpenStudy (jack1):

hahaha! after all that effort ;(

OpenStudy (anonymous):

I'm sorry. Internet difficulties.

OpenStudy (loser66):

to h(g(x)). At h (x) wherever you see x, replace by g(x) (the whole g(x)) and then simplify it. Do the same with g(h(x)) . Your problem: